# OriGen: Enhancing RTL Code Generation with Code-to-Code Augmentation and Self-Reflection

### v1

@@ 3.2. Code-to-Code Augmentation
3.3. Error-Correction Dataset
Given the observed performance gap, we decide to adopt the knowledge distillation method to acquire stronger error understanding and self-reflection capabilities from the closed-source LLM. During the code-to-code augmentation process discussed in Section 3.2, the closed-source LLM generates a diverse set of RTL code samples that fail to compile, which are then corrected in the subsequent self-reflection process. We select the code samples generated in the code-to-code augmentation process that pass compilation after correction, along with the corresponding samples that failed to compile. The natural language instruction descriptions and compiler error messages associated with these samples serve as data sources for constructing the error-correction dataset. This approach enables the model to be exposed to a broad spectrum of error types and their respective corrections, facilitating the development of robust error understanding and self-reflection capabilities.
3.4. Code Generation and Fix
@@
The generative and iterative self-reflection process is illustrated in Figure 2. Initially, OriGen accepts the user’s natural language specification and proceeds to generate the corresponding RTL code using Gen LoRA. The generated code is then subjected to the compiler’s verification process using Iverilog. In cases where the code fails to compile, OriGen initiates a self-reflection process to fix the code. This process involves utilizing the erroneous code, compiler error messages, and natural language instructions to guide the model in identifying and rectifying the issues. The self-reflection loop continues until the generated code successfully passes compilation or reaches a predefined maximum number of iterations.
An example of self-reflection is illustrated in Figure 5. The generated code initially contains a syntactic error where the wires ply_temp and p2y_temp are assigned values within always blocks, which is not permitted in Verilog as wires cannot serve as left-hand side values for procedural assignments. This error is subsequently rectified by leveraging the Fix LoRA model during the self-reflection process, using the instruction template shown in Figure 6. The rectified code employs the assign statement to perform the assignments, rather than attempting to assign values within the always blocks, which was the source of the initial error.
3.5. VerilogFixEval
To evaluate the capability of different models in reflecting on and improving from Verilog compiler error messages, we constructed the VerilogFixEval benchmark. The benchmark comprises code samples that failed to pass the compilation verification, along with the corresponding natural language instructions and compiler error messages. The erroneous RTL code samples were selected from code generated by LLMs that performed comparable with GPT-3.5 on the VerilogEval benchmark. This selection approach was adopted since OriGen achieves a relatively high compilation pass rate. By including code samples from poorly performing LLMs, the benchmark ensures the diversity of errors while avoiding potential bias in the test results that may favor OriGen due to errors generated by OriGen itself.
During the evaluation process, the model will receive natural language instructions, erroneous RTL code, and compiler error messages, and will be required to correct the RTL code. The final evaluation metrics include two parts, syntactic correctness and functional correctness.
@@ 4.1. Experimental Setting
For our pre-trained model, we selected the DeepSeek-Coder-7B-Instruct model, as it exhibits the best performance in Verilog code generation among all 7B models, to the best of our knowledge. It is to ensure that our pre-trained model possesses strong capabilities in the domain of Verilog code generation, providing a solid foundation for further fine-tuning and evaluation.
To evaluate the performance of Verilog code generation, we selected two representative benchmarks: VerilogEval (Liu et al., 2023c) and RTLLM (Lu et al., 2024). The former, VerilogEval, originates from approximately 150 Verilog tasks on the HDLBits website, which were manually converted to create VerilogEval-Human and generated by GPT-3.5 to produce VerilogEval-Machine. The latter benchmark, RTLLM, consists of 29 Verilog tasks with more diverse levels of difficulty, closely aligned with real-world design tasks. Both benchmarks employ the widely adopted pass@k evaluation metric to assess the correctness of the generated code’s functionality. In this metric, if any one of the k samples passes the unit test, the problem is considered solved.
4.2. Model Training
@@
[1.2pt]
We employ the LoRA (Low-Rank Adaptation)  (Hu et al., 2021) method to train the model’s capability in generating RTL code. This approach allows for enhancing the model’s specific abilities in RTL code generation while minimizing the impact on its other capabilities. For all the training processes, we employ the float16 mixed precision method, although the model is trained in bfloat16 precision. We utilize the Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and the cosine learning rate decay to schedule our learning rate. The warm-up ratio is set to 0.03.
During the training process, we initially segment the dataset into smaller subsets, each containing 10k samples, resulting in a total of 18 groups. We then perform fine-tuning on these smaller datasets to evaluate their quality and assess the model’s performance training on them. Throughout the experiments, it is evident that the dataset’s performance is relatively sensitive to the learning rate.
As illustrated in Figure 7, we choose several subsets of the dataset, including 6 smaller 10k-samples datasets and 2 larger 30k-samples datasets. Employing varying learning rates during the training phase lead to noticeable differences in the pass@1 metric on the VerilogEval benchmark  (Liu et al., 2023c). Simultaneously, the results obtained from training on different datasets exhibit significant variations. These observations highlight the sensitivity of the model’s performance to the choice of learning rate and the characteristics of the training dataset. In response, we implement a strategy of training each small dataset with its optimal learning rate. Furthermore, we conduct training exclusively on those datasets where the pass@1 metric surpass 42%. This approach mitigate the sensitivity of the model’s performance to the learning rate, ensuring that each dataset subset is trained with the most suitable learning rate to achieve optimal results. Additionally, it avoids the detrimental influence of underperforming datasets.
Figure 7. Impact of Learning Rate for Different Datasets
@@ 4.3. Functional Correctness
In the VerilogEval benchmark  (Liu et al., 2023c), for pass@1 in Human and Machine categories, OriGen achieves 76.1% and 51.4% respectively. It outperforms remarkably all other alternatives designed for RTL code generation, for example, compared to the best-performed fully open-source RTLCoder (Liu et al., 2023b), OriGen surpasses it by 9.8% in Human benchmark. When compared with commercial closed-source models, OriGen also achieves excellent performance. It significantly surpasses GPT-4 of 2023 version, which was previously the widely compared baseline. Even when compared with the state-of-the-art models, OriGen outperforms Claude3-Haiku/Sonnet, only slightly inferior to the current best models GPT-4 Turbo and Claude3-Opus by 3% in Human benchmark. Moreover, as evident from Table 4.2, it achieves the best performance on the Machine benchmark, remarkably outperforming other models including GPT-4 Turbo.
To prevent our model from over-fitting on the VerilogEval benchmark, we also conduct experiments on another benchmark designed for RTL code generation, RTLLM  (Lu et al., 2024). Similar results are observed on RTLLM as shown in Table 4.2, where OriGen significantly outperforms other models, achieving performance comparable to GPT-4 Turbo and Claude3-Opus.
In summary, OriGen outperforms all non-commercial models remarkably in all metrics on both benchmarks.
Moreover, as illustrated in Figure 8, we observe that the strategy of training each small dataset with its optimal learning rate and excluding the datasets with relatively poor performance allows the model’s performance to gradually improve as the dataset grows larger. This demonstrates the effectiveness of the training methodology.
@@ 4.4. Capability of Self-Reflection
As illustrated in Table 4.4, experimental results on VerilogFixEval demonstrate that OriGen achieves the best performance in syntactic error correction, surpassing GPT-4 Turbo by 18.1%. This indicates that after training on the code-correction dataset, the model acquires powerful self-reflection and code rectification capabilities. For functional correctness, OriGen significantly outperforms RTLCoder by 24.5% and falls short of the best-performing model, Claude3-Opus, by only 6.8%. This showcases OriGen’s ability to consider the functional correctness of the generated code while rectifying syntactic errors.
@@ 4.5. Ablation Studies
We perform two ablation experiments to investigate the efficacy of the code-to-code augmentation method and to examine the model’s self-reflection capability before and after training on the error-correction dataset.
For the ablation study of the code-to-code data augmentation method, we compared the performance of models trained on RTL code dataset before and after code-to-code augmentation method.
@@
VerilogEval-Human (Liu et al., 2023c) Pass@1 38.7 51.4
VerilogEval-Machine (Liu et al., 2023c) Pass@1 62.5 76.2
@@
baseline 53.4 24.4
baseline + error message 62.9 25.8
finetune 82.3 31.6
finetune + error message 87.3 33.0
@@
This paper introduces OriGen, an open-source framework for RTL code generation. It proposes a novel code-to-code augmentation methodology to generate high-quality, large-scale RTL code datasets, which enhances the model’s training data. The framework also introduces a self-reflection mechanism that allows OriGen to autonomously fix syntactic errors by leveraging compiler feedback, thereby improving its code generation accuracy. Furthermore, we construct a dataset to improve the model’s capability of self-reflection based on compiler error messages and erroneous code and develop a benchmark to evaluate this capability. Experimental results demonstrate that OriGen remarkably outperforms other open-source alternatives in RTL code generation, surpassing the previous best-performing LLM by 9.8% on the VerilogEval-Human benchmark and is comparable with GPT-4. Moreover, OriGen exhibits superior capabilities in self-reflection and error rectification, surpassing GPT-4 by 18.1% in syntactic correctness on the VerilogFixEval benchmark.

### v2

@@ 3.4. Code Generation and Fix
Figure 5 illustrates an example of self-reflection in the correction of Verilog code for a module that implements the functionality of the 7458 chip, which includes four AND gates and two OR gates.
The generated code initially contains a syntactic error: wires ‘ply_temp‘ and ‘p2y_temp‘ are incorrectly assigned values within always blocks. In Verilog, wires cannot be assigned values on the left-hand side of procedural assignments. This error is corrected by using the Fix LoRA model during the self-reflection process, following the instruction template shown in Figure 6. The regenerated code resolves these issues by employing continuous assignments for ‘p1y‘ and ‘p2y‘, rather than assigning values within always blocks, ensuring that the logic accurately reflects the intended operations.
3.5. VerilogFixEval Benchmark
To assess the ability of various models to reflect and improve from Verilog compiler error messages, we developed the VerilogFixEval benchmark. This benchmark consists of code samples that failed compilation verification, along with corresponding natural language instructions and compiler error messages. The faulty RTL code samples were selected from those generated by LLMs that performed similarly to GPT-3.5 on the VerilogEval benchmark. This selection method was adopted because OriGen achieves a relatively high compilation pass rate. By including code samples from poorly performing LLMs, the benchmark ensures a diversity of errors while minimizing potential bias in test results that could favor OriGen due to errors produced by OriGen itself.
During the evaluation process, the model will receive natural language instructions, faulty RTL code, and compiler error messages to correct the RTL code. The final evaluation metrics consist of two components: syntactic correctness and functional correctness.
@@ 4.1. Experimental Setting
[1.2pt] For our pre-trained model, we selected the DeepSeek-Coder-7B-Instruct model, as it exhibits the best performance in Verilog code generation among all 7B models, to the best of our knowledge. It is to ensure that our pre-trained model possesses strong capabilities in the domain of Verilog code generation, providing a solid foundation for further fine-tuning and evaluation.
To evaluate the performance of Verilog code generation, we selected two representative benchmarks: VerilogEval (Liu et al., 2023c) and RTLLM (Lu et al., 2024). The former, VerilogEval, originated from approximately 150 Verilog tasks on the HDLBits website, which were manually converted to create VerilogEval-Human and generated by GPT-3.5 to produce VerilogEval-Machine. The latter benchmark, RTLLM, consists of 29 Verilog tasks with more diverse levels of difficulty, closely aligned with real-world design tasks. Both benchmarks employ the widely adopted pass@k evaluation metric to assess the correctness of the generated code’s functionality. In this metric, if any one of the k samples passes the unit test, the problem is considered solved.
4.2. Model Training
We employ the LoRA (Low-Rank Adaptation)  (Hu et al., 2021) method to train the model’s capability in generating RTL code. This approach allows for enhancing the model’s specific abilities in RTL code generation while minimizing the impact on its other capabilities. For all the training processes, we employ the float16 mixed precision method, although the model is trained in bfloat16 precision. We use the AdamW optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$, along with cosine learning rate decay for scheduling. The warm-up ratio is set to 0.03, and the batch size is 8.
@@ 4.3. Functional Correctness
In the VerilogEval benchmark  (Liu et al., 2023c), OriGen achieves pass@1 scores of 54.4% in the Human category and 74.1% in the Machine category, outperforming remarkably all other alternatives designed for RTL code generation. For instance, compared to the best-performing fully open-source RTLCoder (Liu et al., 2023b), OriGen surpasses it by 12.8% on the Human benchmark. Additionally, compared to commercial closed-source models, OriGen demonstrates exceptional performance. It significantly outperforms GPT-4 (2023), which was previously considered a key benchmark. Furthermore, it exceeds the teacher model Claude3-Haiku, underscoring the effectiveness of filtering and self-reflection in code augmentation.
Among the state-of-the-art models, OriGen outperforms Claude3-Haiku/Sonnet across all metrics and surpasses GPT-4 Turbo in the pass@1 metric while being only slightly inferior to Claude3-Opus by less than 1% on the Human benchmark pass@1 metric. Moreover, as shown in Table 4.1, it achieves the best performance on the Machine benchmark, notably outperforming other models including Claude3-Opus and GPT-4 Turbo. The performance gap between Human and Machine categories primarily stems from descriptions generated by GPT-3.5 for the Machine category. In contrast, OriGen is trained on a vast amount of synthesized data, enabling it to excel with problems generated by LLMs.
@@
We also conduct experiments to investigate the relationship between dataset size and model performance, as depicted in Figure7. The performance is evaluated using the pass@1 metric on VerilogEval Human. We examined dataset sizes ranging from 0 (representing the base model without fine-tuning) to approximately 180,000 samples (utilizing the entire fine-tuning dataset).
The results demonstrate a clear positive correlation between dataset size and model performance. Starting from a pass@1 score of 31.7 for the base model, we observe a sharp initial increase to 43.2 with just 10k samples. This is followed by a more gradual but consistent improvement as the dataset size increases, eventually reaching a peak performance of 54.4 with the full dataset of 180k samples. Notably, the performance gains are not linear. The most substantial improvements occur in the early stages of data addition, with diminishing returns as the dataset size approaches its maximum. This pattern underscores the critical role of data quantity and diversity in model performance for further improvements beyond a certain scale.
To prevent our model from over-fitting on the VerilogEval benchmark, another benchmark RTLLM  (Lu et al., 2024) is also used to evaluate. Similar results are observed on RTLLM as shown in Table 4.1, where OriGen significantly outperforms other models, achieving performance comparable to GPT-4 Turbo and Claude3-Opus.
In summary, OriGen significantly outperforms all non-commercial models across all metrics on both benchmarks and achieves comparable performance to the current state-of-the-art commercial closed-source models.
@@ 4.4. Capability of Self-Reflection
As illustrated in Table 4.4, experimental results on VerilogFixEval demonstrate that OriGen achieves the best performance in syntactic error correction, surpassing GPT-4 Turbo by 19.9%. This indicates that after training on the code-correction dataset, the model acquires powerful self-reflection and code rectification capabilities. For functional correctness, OriGen significantly outperforms RTLCoder by 18.2% and falls short of the best-performing model, Claude3-Opus, by only 6.3%. This showcases OriGen’s ability to consider the functional correctness of the generated code while rectifying syntactic errors.
@@ 4.5. Ablation Studies
We perform two ablation experiments to investigate the efficacy of the code-to-code augmentation method and to examine the model’s self-reflection capability before and after training on the error-correction dataset. For the ablation study of the code-to-code data augmentation method, we compared the performance of models trained on RTL code dataset before and after code-to-code augmentation method.
@@
VerilogEval-Human (Liu et al., 2023c) Pass@1 41.6 54.4
VerilogEval-Machine (Liu et al., 2023c) Pass@1 62.5 74.1
@@
baseline 53.8 23.4
baseline + error message 63.5 25.6
finetune 82.7 31.9
finetune + error message 89.1 33.5
@@
This paper introduces OriGen, an open-source framework for RTL code generation. It proposes a novel code-to-code augmentation methodology to generate high-quality, large-scale RTL code datasets, which enhances the model’s training data. The framework also introduces a self-reflection mechanism that allows OriGen to autonomously fix syntactic errors by leveraging compiler feedback, thereby improving its code generation accuracy. Furthermore, we construct a dataset to improve the model’s capability of self-reflection based on compiler error messages and erroneous code and develop a benchmark to evaluate this capability. Experimental results demonstrate that OriGen remarkably outperforms other open-source alternatives in RTL code generation, surpassing the previous best-performing LLM by 12.8% on the VerilogEval-Human benchmark and is comparable with GPT-4 Turbo. Moreover, OriGen exhibits superior capabilities in self-reflection and error rectification, surpassing GPT-4 by 19.9% in syntactic correctness on the VerilogFixEval benchmark.